Write an equation for the set of all points in the plane equidistant from (0,–6) and y= –10.

Write your answer in vertex form. Simplify any fractions.


1
Expert's answer
2021-12-22T04:11:10-0500

Let the parametric point be (x,y) on the plane which is equidistant from the given point (0,6)  hence, we have

"\sqrt{x^2+(y+6)^2}=y+10\\
=x^2+y^2+12y+36=(y+10)^2\\
= x^2+y^2+12y+36=y^2+20y+100\\
=x^2-8y-64=0"